# Wigner’s friend scenarios: on what to condition and how to verify the predictions

### v1

@@ Abstract
Wigner’s friend experiment and its modern extensions display the ambiguity of the quantum mechanical description regarding the assignment of quantum states. While the friend applies the state-update rule to the system upon observing an outcome of her measurement in a quantum system, Wigner describes the friend’s measurement as a unitary evolution, resulting in an entangled state for the composite system of the friend and the system. In this respect, Wigner is often referred to as a “superobserver” who has the supreme technological ability to keep the friend’s laboratory coherent. As such, it is often argued that he has the “correct” description of the state. Here we show that the situation is more symmetrical than is usually thought: there are different types of information that each of the observers has that the other fundamentally cannot have - they reside in different “bubbles” (in Calvalcanti’s terminology). While this can explain why the objectivity of the state assignment is only relative to the bubble, we consider more elaborated situations in form of a game in which the players can switch between bubbles. We find that, in certain circumstances, observers may be entitled to adopt and verify the state assignment from another bubble if they condition their predictions on all information that is in principle available to them.
@@ I Introduction
The notorious quantum measurement problem stems from an ambiguity about two alternative dynamics that are present in quantum theory. In the words of Baumann and Wolf [1], “in what is called standard quantum mechanics, the measurement-update rule, commonly associated with a collapse, is a break with the otherwise unitary evolution governed by the Schrödinger equation. The formalism, however, provides no indication about when to apply this rule: It does not state what qualifies some interactions as measurements but not others (the measurement problem).”
@@
The problem is most notably exemplified by the Wigner’s friend thought experiment [2], whose scope has been enlarged in recent years [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. This takes the ambiguity about which dynamical law to apply to the extreme by featuring an observer –historically referred to as the friend– who takes both the roles of an observer performing a quantum experiment and of a quantum system observed by another observer, Wigner. More specifically, the friend performs a measurement on a quantum system in her laboratory, while Wigner is able to perform quantum measurement on the whole laboratory, thus treating both the friend and the particle as a joint quantum system.
@@
What is the state of the system after the friend’s measurement? Since the friend observes a definite outcome, it seems legit to say that the state should be updated accordingly, thus applying the “collapse” rule. On the other hand, from Wigner’s perspective, the friend and the system are merely two interacting subsystems of an isolated system, thus he seems entitled to describe the same physical process unitarily. The question then arises: is there a single, absolute, “correct” state? By this we mean that every hypothetical observer ought to assign the same state in order to make predictions that turn out to be empirically correct.
@@
The answer in the positive view has been recently challenged by a series of no-go theorems [3, 5, 4, 11, 12, 19, 15, 18, 20], in so far as it is possible that there exist no “single reality” that can be shared between the friend and Wigner, and they instead live in different “bubbles” [12].  [^1] This is reminiscent of the philosophical position of fragmentalism [23, 24], which upholds “that the world is not a monolithic whole constructed from mutually compatible facts, but rather a collection of fragments, with each fragment containing mutually compatible facts, while different fragments are incompatible” [25]. A bubble is an information-theoretic concept and is to be understood as a “region” in which certain information about, e.g., a measurement result, is (in principle) available. Every hypothetical observer who has in principle access to the same relevant classical information –i.e., that can be copied, broadcasted, etc.– lives in the same bubble. In formal terms, these no-go theorems state that, under certain assumptions, there is no joint probability for the results obtained by the friend and Wigner such that the respective probabilities for their results can be understood as marginals of it. This resolves the ambiguity of the state assignment, insofar as a state can be considered objective only relatively to a specified bubble, i.e., each and every hypothetical observer within the same bubble will be entitled to assign the same state within their respective bubble. However, the states will in general be different for different bubbles.
@@
The strategies for rational prediction are less clear when the observers can “switch” bubbles, for example, when a measurement is performed in another bubble, but the player finds a way to verify the outcomes of this measurement and thus their previous predictions. In such cases, if an observer who resides in one bubble has access to a quantum state assignment from another bubble, should they adopt it for the purpose of predicting the probabilities for the occurrence of the measurement outcomes? Imagine, for example, that the friend knows the state in which Wigner describes her laboratory, and she also knows which measurement of her laboratory Wigner will make next. On whose knowledge should she condition her prediction of the result of this measurement, her own or Wigner’s? Given that the measurement will affect her memory, will she even be able to evaluate her previous prediction after the measurement has been completed?
In this paper we address these questions by considering a game played between the friend and Wigner. We will argue that, in certain circumstances, both players can be persuaded that their state assignment, based solely on the information available in their individual bubbles, is not optimal for predicting future measurements, but that of the other player is. In some cases, the players will even be entitled to use the information that is available in other bubbles to make predictions about the measurements made in those bubbles, in addition to their own perspective when making predictions about measurements made in their own bubble. To this end, we discuss to what extent one can assume the same identity of an observer playing the game, i.e. making predictions, undergoing measurements of their own bubble including their memories, and verifying their previous predictions.
@@
The first step towards investigating situation when an observer may “switch” the bubbles was taken in Ref. [8], but only for the case where the friend was placed in the situation. Here we aim at the situations more symmetrical for the two observers,. Hence, we also analyse the game when the knowledge of the two observers is not only maximally incompatible (i.e. when the two measurements in the respective laboratories are maximally non-commutative), but partially overlapping, thus covering the whole spectrum from full compatibility to maximal incompatibility. In that sense our work amounts to the task of finding information-theoretic conditions that allow to univocally decide which state assignment is rational to adopt in cases when an observer can “switch” from one bubble to another and when some information is allowed to be exchanged between bubbles.
II Interacting “bubbles”
@@
Let us consider the qubit system (S) and the friend’s (F) and Wigner’s (W) environments or laboratories. At the beginning the situations of the friend and of Wigner are fully symmetrical and similar to the one experienced by our experimental colleagues in everyday practice. In particular, we assume that they each possess reliable experimental devices and act as rational agents. Moreover, the interaction between the quantum system and the environment in each laboratory results in environmentally-induced decoherence and a preferred environmental basis in which measurement outcomes are redundantly recorded [26, 27, 28, 29]. The records may include “click” in a detector, laboratory notes or even the observer’s memory. Nevertheless, for the purposes of our argument, we will assume that the preferred bases of the friend’s and Wigner’s environments are not identical, and that there is a “boundary region” between two bubbles, in which a unitary transformation maps one basis into the other.
@@
where $|\psi\rangle_{S}$ is a state of the (qubit) system and $|\phi_{0}\rangle_{F}$, $|\omega_{0}\rangle_{W}$ are no further specified “ready” states of the friend’s and Wigner’s environment, rspectively. We assume for simplicity $|\psi\rangle_{S}=\frac{1}{\sqrt{2}}(|\uparrow\rangle_{S}+|\downarrow\rangle_{S})$ where $\{|\uparrow\rangle_{S},|\downarrow\rangle_{S}\}$ is the orthonormal basis of the qubit in which the friend measures, say the $z$-direction of the spin.
@@
On the other hand, the friend performing the measurement on the qubit system enforces the measurement (state-update) postulate by the application of projectors $\Pi_{\uparrow}=|\uparrow\rangle_{S}\langle\uparrow|$ and $\Pi_{\downarrow}=|\downarrow\rangle_{S}\langle\downarrow|$ over $|\psi_{s}\rangle$ in Eq. (1). An equivalent perspective on the measurement process for the friend is to first adopt the state assignment (2) for the qubit and her environment, while concurrently applying the projector $\Pi^{F}_{0}=|\phi_{0}\rangle_{F}\langle\phi_{0}|$ and $\Pi^{F}_{1}=|\phi_{1}\rangle_{F}\langle\phi_{1}|$ over $|\Psi^{W}_{1}\rangle$ in Eq. (2) [^2] Provided that for the friend only reading out the states in the preferred basis of the environment is permitted, two friend’s descriptions either by applying the state update rule to the quantum system in state (1) or by applying the state update rule to his environment in state  (2) are equivalent. Indeed, when performing experiments, our experimental colleagues apply the state update rule to the measured system. Nevertheless, they may still employ description  (2) to analyse the decoherence process induced by the measurement. Note that in either case, the state of the Wigner’s environment is irrelevant and can be traced out.. This leads to classically correlating the friend environment to the measurement result in the following states, depending on the friend’s measurement outcome:
@@
Let us consider now that a hypothetical physical process at the boundary of the two labs couples the two environments as illustrated in Fig. 1 and both observers knows the unitary. This coupling is described by a unitary operator $U_{\mathrm{int}}$ acting over the system and the two environments, defined such that the only elements on which it acts non-trivially are:
@@
where $|\Phi^{\pm}\rangle_{SF}=(|\uparrow\rangle_{S}|\phi_{0}\rangle_{F}\pm|\downarrow\rangle_{S}|\phi_{1}\rangle_{F})/\sqrt{2}$ are two of the standard Bell’s states over the system’s and friend’s Hilbert space. Therefore, by applying $U_{\text{int}}$ to Eq. (2) we obtain for Wigner’s description:
@@
The additional unitary operation does not affect the initial correlations between the state of the qubit and the friend’s environment. Hence, looking at his environmental degree of freedom, Wigner can only access the information about the system and the friend being perfectly correlated in state $|\Phi^{+}\rangle$, while the actual information about “which-outcome” $\uparrow$ or $\downarrow$ of the qubit measurement is inaccessible. In Appendix A we consider more general unitaries where partial information about friend’s measurement outcome is leaked into Wigner’s environment.
@@
associated respectively to $\phi_{0}$ and $\phi_{1}$ results. The above state can be equivalently written as
@@
with $|\omega_{\pm}\rangle_{W}=(|\omega_{0}\rangle_{W}\pm|\omega_{1}\rangle_{W})/\sqrt{2}$. From Eq. (6), it can be seen that Wigner’s environmental states $\{|\omega_{0}\rangle_{W},|\omega_{1}\rangle_{W}$ are perfectly correlated with the Bell states of the system and the friend’s environment. This implies that no information about the outcome of the friend’s measurement can be inferred from the preferred environmental basis in Wigner’s laboratory. Moreover, from Eq. (7), it can be seen that for any of the two outcomes of the qubit measurement and following the action of $U_{\mathrm{int}}$, the qubit system and the friend’s laboratory become entangled in a state with two “branches”. In each branch, the qubit state and the friend’s laboratory records are found to be perfectly correlated for any measurement result: if the system is $\uparrow$, the state of the record is $\phi_{0}$, and if it is $\downarrow$, the record is in $\phi_{1}$ (and it is never so that e.g. outcome $\uparrow$ is correlated with $\phi_{1}$). This demonstrates that the correlation between the spin and the laboratory’s record is still maintained after the unitary interaction. We will analyze the meaning of this state in greater detail in Section V.
For the purpose of the game, we will assume that after the first measurement of the qubit by the friend, one of two possible measurements will be performed next. Either the friend will repeat the same measurement ($\text{M}_{F}$) of her qubit, measuring again the spin along the $z$ direction, or alternatively, Wigner will perform a measurement (MW) of the entire friend’s laboratory in the preferred basis red of his environment, i.e., using projectors $\Pi_{0}^{W}=|\omega_{0}\rangle\langle\omega_{0}|_{W}$ and $\Pi_{1}^{W}=|\omega_{1}\rangle\langle\omega_{1}|_{W}$. In the game, the two observers will be required to provide their predictions regarding the outcomes of one or the other measurement.
III The rules of the game
Next we will introduce the rules of the game. The game involves two players, Wigner (W) and the friend (F), as well as two referees, RW and RF, located in Wigner’s and the friend’s bubbles, respectively.
@@ III The rules of the game
It is important to note that all forms of communication, including the submission of predictions and the sharing of random choices regarding the next measurement, will not contain any information about the outcome of the first friend’s measurement and will not affect the states as assigned in the bubbles. In this sense, a bubble is an information theoretic concept, rather defined as the kind of information in principle available than to an amount of leaked information (i.e., there is no need of complete isolation between two regions for them to be two bubbles).
We will now proceed to comment on the rules of the game. The players need not rely on the honesty of the referees in step 5. in evaluating their predictions. Instead, they can inspect their previous predictions in form of permanent records (e.g., signed pieces of paper on which their prediction is written) that they handed to the responsible referee. Also, for case $M_{W}$ the friend’s memory will be subjected to a measurement thereby erasing that (tiny) part of their memory which recorded the outcome of the previous measurement. This may give rise to the concern that the “forgetful” observer may not agree that they ever made any prediction, or at least not the one which they will be faced with at the end of the game. However, since the players may still be presented with a record of their previous predictions –the signed piece of paper with their signature– this may provide them with sufficient evidence that the prediction was indeed made by them.
Finally, in order to avoid potential difficulties arising from the so-called self-reference problem [30], namely, the circumstance that the friend might have to measure her own state, let us append to the already considered scenario a further auxiliary memory system (A) of the same dimension than $F$, prepared in an initial “ready” state $|R\rangle_{A}$. Applying then a unitary swap operation, $U_{\mathrm{SWAP}}$ acting over friend and the auxiliary system, the memory of the friend is coherently transferred to the auxiliary system (such that it swaps the indices of the friend and the auxiliary system). Under the action of $U_{\mathrm{SWAP}}$, the states in Eq. (5) and (6) transforms as
@@
IV Measurement in Wigner’s bubble

### v2

@@ Abstract
Wigner’s friend experiment and its modern extensions display the ambiguity of the quantum mechanical description regarding the assignment of quantum states. While the friend applies the state-update rule to the system upon observing an outcome of her measurement in a quantum system, Wigner describes the friend’s measurement as a unitary evolution, resulting in an entangled state for the composite system of the friend and the system. In this respect, Wigner is often referred to as a “superobserver” who has the supreme technological ability to keep the friend’s laboratory coherent. As such, it is often argued that he has the “correct” description of the state. Here we show that the situation is more symmetrical than is usually thought: there are different types of information that each of the observers has that the other fundamentally cannot have - they reside in different “bubbles” (in Calvalcanti’s terminology). While this can explain why the objectivity of the state assignment is only relative to the bubble, we consider more elaborated situations in the form of a game in which the players can switch between bubbles. We find that, in certain circumstances, observers may be entitled to adopt and verify the state assignment from another bubble if they condition their predictions on all information that is in principle available to them.
@@ I Introduction
The notorious quantum measurement problem stems from an ambiguity between two alternative dynamics that are present in quantum theory. In the words of Baumann and Wolf [1], “in what is called standard quantum mechanics, the measurement-update rule, commonly associated with a collapse, is a break with the otherwise unitary evolution governed by the Schrödinger equation. The formalism, however, provides no indication about when to apply this rule: It does not state what qualifies some interactions as measurements but not others (the measurement problem).”
@@
The problem is most notably exemplified by the Wigner’s friend thought experiment [2], whose scope has been enlarged in recent years [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. This takes the ambiguity about which dynamical law to apply to the extreme by featuring an observer—historically referred to as the friend—who takes both the roles of an observer performing a quantum experiment and of a quantum system observed by another observer, Wigner. More specifically, the friend performs a measurement on a quantum system in her laboratory, while Wigner is able to perform quantum measurement on the whole laboratory, thus treating both the friend and the particle as a joint quantum system.
@@
What is the state of the system after the friend’s measurement? Since the friend observes a definite outcome, it seems legit to say that the state should be updated accordingly, thus applying the “collapse” rule. On the other hand, from Wigner’s perspective, the friend and the system are merely two interacting subsystems of an isolated system, thus he seems entitled to describe the same physical process unitarily. The question then arises: is there a single, absolute, “correct” state? By this, we mean that every hypothetical observer should assign the same state, which can then be used to make empirically verifiable predictions consistent with quantum theory.
@@
The answer in the positive view has been recently challenged by a series of no-go theorems [3, 5, 4, 11, 12, 19, 15, 18, 20], in so far as it is possible that there exist no “single reality” that can be shared between the friend and Wigner. It was proposed instead that Wigner and his friend could live in different, incommensurable realities, called “bubbles” [12]. [^1] The concept of bubble, as it will be used here, is very similar to Healey’s “agent situation” [24]. These ideas are reminiscent of the philosophical position of fragmentalism [25, 26], which upholds “that the world is not a monolithic whole constructed from mutually compatible facts, but rather a collection of fragments, with each fragment containing mutually compatible facts, while different fragments are incompatible” [27].
In this work, we (i) refine and further formalize the concept of a “bubble” as an information-theoretic notion, and model communication between bubbles; (ii) devise a scenario (in the form of a gambling game) in which an observer living in one bubble can be persuaded that their state assignment—based solely on the information available from measurement results obtained within their own bubble—is not optimal for predicting future measurements, whereas the state assignment of an observer in another bubble is; and (iii) demonstrate, for the first time, that the friend—typically regarded as less powerful than the superobserver Wigner—can make more precise predictions than the latter, thereby symmetrizing the narrative.
II “Information Bubbles”
To describe the incommesurability of the measurement outcomes by different observers in Wigner’s friend extended scenarios, Cavalcanti introduced the concept of “bubbles” [12]. Building on this intuition, a bubble can be defined as the ”locus” where the same relevant information—for example, a measurement result—is, in principle, available. Every hypothetical observer who has in principle access to the same relevant classical information—i.e., that can be copied, broadcasted, etc.—lives, therefore, in the same bubble. To this extent, one usually takes Wigner to live in a different bubble than the friend because the latter is enclosed in a fully isolated laboratory. Yet, a bubble is notably an information-theoretic concept and, as such, is fundamentally independent of space-time location [28], relying exclusively on the relevant information that is, in principle, retrievable. Thus, in general, the existence of the bubble is not compromised even if information is exchanged between the parties, provided that this information is uncorrelated with the outcomes of the quantum measurements. (An asymmetry in the flow of outcome-related information from the friend to Wigner, as opposed to from Wigner to the friend, will be discussed later.)
The already mentioned no-go theorems for extended Wigner’s friend scenarios [3, 5, 4, 11] formally state that, under certain assumptions, there is no joint probability for the results obtained by the friend and Wigner such that the respective probabilities for their results can be understood as marginals of it. The bubble view is a way to resolve the ambiguity of the state assignment, insofar as a state can be considered objective only relatively to a specified bubble, i.e., each and every hypothetical observer within the same bubble will be entitled to assign the same state within their respective bubble. However, the states will in general be different for different bubbles.
@@
The strategies for rational predictions are less clear when the observers can “switch” bubbles, for example, when a measurement is performed in another bubble, but the player finds a way to verify the outcomes of this measurement and thus their previous predictions. In such cases, if an observer who resides in one bubble has access to a quantum state assignment from another bubble, should they adopt it for the purpose of predicting the probabilities for the occurrence of the measurement outcomes? Imagine, for example, that the friend knows the state in which Wigner describes her laboratory, and she also knows which measurement of her laboratory Wigner will make next. Should she condition her predictions solely on the knowledge of the measurement results she observed in her lab, or on Wigner’s knowledge. And given that the future measurement will affect her memory, will she even be able to evaluate her previous prediction after the measurement has been completed?
@@
The first step towards investigating situation when an observer may “switch” the bubbles was taken in Ref. [8], but only for the case where the friend was placed in the situation. Here we aim at the situations more symmetrical for the two observers: we propose a game (in Sect. IV.1) when the knowledge of the two observers is not only maximally incompatible (i.e. when the two measurements in the respective laboratories are maximally non-commutative), but partially overlapping, thus covering the whole spectrum from full compatibility to maximal incompatibility. In that sense our work amounts to the task of finding information-theoretic conditions that allow to univocally decide which state assignment is rational to adopt in cases when an observer can “switch” from one bubble to another and when some information is allowed to be exchanged between bubbles.
III Communicating “bubbles”
@@
Let us consider the qubit system (S) and the friend’s (F) and Wigner’s (W) environments or laboratories. At the beginning the situations of the friend and of Wigner are fully symmetrical and similar to the one experienced by our experimental colleagues in everyday practice. In particular, we assume that they each possess reliable experimental devices and act as rational agents. Moreover, the interaction between the quantum system and the environment in each laboratory results in environmentally-induced decoherence and a preferred environmental basis in which measurement outcomes are redundantly recorded [29, 30, 31, 32]. The records may include “clicks” in a detector, laboratory notes or even the observer’s memory. Nevertheless, for the purposes of our argument, we will assume that the preferred bases of the friend’s and Wigner’s environments are not identical, and that there is a “boundary” between two bubbles, in which a unitary transformation maps one basis into the other.
@@
where $|\psi\rangle_{S}$ is a state of the (qubit) system and $|\phi_{0}\rangle_{F}$, $|\omega_{0}\rangle_{W}$ are no further specified “ready” states of the friend’s and Wigner’s environment, respectively. At this stage, the state assignment could be one upon which both the friend and Wigner agree. We assume for simplicity $|\psi\rangle_{S}=\frac{1}{\sqrt{2}}(|\uparrow\rangle_{S}+|\downarrow\rangle_{S})$ where $\{|\uparrow\rangle_{S},|\downarrow\rangle_{S}\}$ is the orthonormal basis of the qubit in which the friend measures, say the $z$-direction of the spin.
@@
On the other hand, the friend performing the measurement on the qubit system enforces the measurement (state-update) postulate by the application of projectors $\Pi_{\uparrow}=|\uparrow\rangle_{S}\langle\uparrow|$ and $\Pi_{\downarrow}=|\downarrow\rangle_{S}\langle\downarrow|$ over $|\psi_{s}\rangle$ in Eq. (1). An equivalent perspective on the measurement process for the friend is to first adopt the state assignment (2) for the qubit and her environment, while concurrently applying the projector $\Pi^{F}_{0}=|\phi_{0}\rangle_{F}\langle\phi_{0}|$ and $\Pi^{F}_{1}=|\phi_{1}\rangle_{F}\langle\phi_{1}|$ over $|\Psi^{W}_{1}\rangle$ in Eq. (2). [^2] Provided that the friend can only read out the states in the preferred basis of the environment, two friend’s descriptions are possible. Either by applying the state update rule to the quantum system in state (1) or by applying the state update rule to his environment in state  (2) are equivalent. Indeed, when performing experiments, our experimental colleagues apply the state update rule to the measured system. Nevertheless, they may still employ description  (2) to analyse the decoherence process induced by the measurement. Note that in either case, the state of the Wigner’s environment is irrelevant and can be traced out. This leads to classically correlating the friend environment to the measurement result in the following states, depending on the friend’s measurement outcome:
@@
Let us consider now that a hypothetical physical process at the boundary of the two labs couples the two environments as illustrated in Fig. 1. This coupling is described by a unitary operator $U_{\mathrm{int}}$ acting over the system and the two environments and both observers know this unitary. The unitary is defined through its action on the only states on which it acts non-trivially:
@@
where $|\Phi^{\pm}\rangle_{SF}=(|\uparrow\rangle_{S}|\phi_{0}\rangle_{F}\pm|\downarrow\rangle_{S}|\phi_{1}\rangle_{F})/\sqrt{2}$ are two of the standard Bell’s states over the system’s and friend’s Hilbert spaces. Therefore, by applying $U_{\text{int}}$ to Eq. (2) we obtain for Wigner’s description:
@@
The unitary operation does not affect the initial correlations between the state of the qubit and the friend’s environment. Hence, looking at his environmental degree of freedom, Wigner can only access the information about the system and the friend being perfectly correlated in state $|\Phi^{+}\rangle$, while the actual information about “which-outcome” $\uparrow$ or $\downarrow$ of the qubit measurement is inaccessible. In Appendix A we consider more general unitaries where partial information about friend’s measurement outcome is leaked into Wigner’s environment.
@@
associated respectively to $\phi_{0}$ and $\phi_{1}$ results in Eq. (3). The above state can be equivalently written as
@@
with $|\omega_{\pm}\rangle_{W}=(|\omega_{0}\rangle_{W}\pm|\omega_{1}\rangle_{W})/\sqrt{2}$. From Eq. (6), it can be seen that Wigner’s environmental states $\{|\omega_{0}\rangle_{W},|\omega_{1}\rangle_{W}$ are perfectly correlated with the Bell states of the system and the friend’s environment. This implies that, even in the friend’s state, no information about the outcome of the friend’s measurement can be inferred from the preferred environmental basis in Wigner’s laboratory. Moreover, from Eq. (7), it can be seen that for any of the two outcomes of the qubit measurement and following the action of $U_{\mathrm{int}}$, the qubit system and the friend’s laboratory become entangled in a state with two “branches”. This demonstrates that the correlation between the spin and the laboratory’s record is still maintained after the unitary interaction.
For the purpose of the game, we will assume that after the first measurement of the qubit by the friend and the action of the unitary at the interface, one of two possible measurements will be performed next. Either the friend will repeat the same measurement ($\text{M}_{F}$) of her qubit, measuring again the spin along the $z$ direction, or alternatively, Wigner will perform a measurement ($M_{W}$) of the entire friend’s laboratory in the preferred basis of his environment, i.e., using projectors $\Pi_{0}^{W}=|\omega_{0}\rangle\langle\omega_{0}|_{W}$ and $\Pi_{1}^{W}=|\omega_{1}\rangle\langle\omega_{1}|_{W}$. In the game, the two observers will be required to provide their predictions regarding the outcomes of one or the other measurement.
IV The “Bubble Switching Game”
IV.1 The rules of the game
Next we will introduce the rules of the game. The game involves two players, Wigner (W) and the friend (F), as well as two referees, [^3] To avoid unnecessary complications in the notation, we treat these referees as co-observers who share access to the relevant outcomes within their bubble. While they are not modeled explicitly, we assume they register and agree with the measurement results of the agents in their respective laboratories. RW and RF, located in Wigner’s and the friend’s bubbles, respectively.
@@ IV.1 The rules of the game
It is important to note that all forms of communication, including the submission of predictions and the sharing of random choices regarding the next measurement, will not contain any information about the outcome of the first friend’s measurement and will not affect the states as assigned in the bubbles. In this sense, a bubble is an information theoretic concept, rather linked to the kind of information in principle available than to an amount of leaked non-relevant information (i.e., there is no need of complete isolation between two regions for them to be two bubbles; see Sect. II).
We will now proceed to comment on the rules of the game. Note that the players do not necessarily need to rely on the honesty of the referees in step 5 when evaluating their predictions. Instead, after each run of the game, they can inspect their previous predictions in form of permanent records (e.g., signed pieces of paper on which their prediction is written) that they handed to the responsible referee in step 3. Also, for case $M_{W}$ the friend’s memory will be subjected to a measurement thereby erasing that (tiny) part of their memory which recorded the outcome of the previous measurement. This may give rise to the concern that the “forgetful” observer may not agree that they ever made any prediction, or at least not the one which they will be faced with at the end of the game. However, since the players may still be presented with a record of their previous predictions –the signed piece of paper with their signature– this may provide them with sufficient evidence that the prediction was indeed made by them.
Finally, in order to avoid potential difficulties arising from the so-called self-reference problem [33], namely, the circumstance that the friend might have to measure her own state, let us append to the already considered scenario a further auxiliary memory system (A) of the same dimension than $F$, prepared in an initial “ready” state $|R\rangle_{A}$. Applying then a unitary swap operation, $U_{\mathrm{SWAP}}$ acting over friend and the auxiliary system, the memory of the friend is coherently transferred to the auxiliary system (such that it swaps the indices of the friend and the auxiliary system). Under the action of $U_{\mathrm{SWAP}}$, the states in Eq. (5) and (6) transform as
@@
IV.2 Measurement in Wigner’s bubble